Bob drove home to work at 30mph. After work the traffic was heavier, and he drove home at 25 mph. His driving time to and from work was 1 hour and 6 minutes. How far does he live from his job
Found 2 solutions by TimothyLamb, mananth:
Answer by TimothyLamb(4379)

x = distance from home to work
a = driving time home to work
b = driving time work to home
---
s = d/t
t = d/s
---
driving time home to work:
a = x/30
---
driving time work to home:
b = x/25
---
total time:
a + b = 1 + 6/60
a + b = 1.1
x/30 + x/25 = 1.1
25x/(25*30) + 30x/(25*30) = 1.1
25x + 30x = 1.1*(25*30)
55x = 1.1*(25*30)
x = 1.1*(25*30)/55
x = 15
---
answer:
x = distance from home to work = 15 miles
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by mananth(16946)

let distance from work place to home be x miles
time while going = x/30
time returning = x/25
total time = 1 hour 6 minutes = 1 6/60 = 11/10 hours
x/30 + x/25 = 11/10
multiply equation by LCM 150
150*x/30 + 150*x/25 = 150*11/10
5x+6x=165
11x=165
/11
x=15
Distance = 15 miles